OpenStudy (anonymous):

Let f(x) = x2-4 and g(x) = x+6. a. Find f(g(x)). b. Find g(f(x)).

OpenStudy (bahrom7893):

f(g(x)) = (x+6)^2-4 g(f(x)) = x^2-4 + 6

OpenStudy (anonymous):

.... -.- *facepalm*

OpenStudy (bahrom7893):

myin, check my ansswer, tnx :)

OpenStudy (bahrom7893):

it can all be simplified btw, mad u can do the simplification

OpenStudy (anonymous):

I got something.. but I'm pretty sure it's wrong...

OpenStudy (anonymous):

what is x+6^2

myininaya (myininaya):

\[f(g(x))=(x+6)^2-4\] this is what bahrom typed there

myininaya (myininaya):

that ^ just means we have a power

OpenStudy (anonymous):

How do i simplify x+6^2 is it mean 2x^2 + 36?

myininaya (myininaya):

no no

myininaya (myininaya):

\[(x+6)^2=(x+6)(x+6)=x(x+6)+6(x+6)=x^2+6x+6x+36=x^2+12x+36\]

myininaya (myininaya):

\[(x+6)^2=x^2+12x+36\]

OpenStudy (anonymous):

Ohhhhh I got it wow